katiicat:

what is 55times 4/3? MATH

Extrinix:

For this you can turn the whole number into its fraction form (\(\sf{\dfrac{55}{1}}\)), then, you can multiply 55 by 4, and 1 by 3 \(\sf{\dfrac{55}{1}\times\dfrac{4}{3}}\) 1, What is 55 times 4? and 2, What is 1 times 3?

katiicat:

ok

Extrinix:

You just put those answers on top of each other and you’re good 👍 \(\sf{\dfrac{55 \times 4}{1 \times 3}}\)